### v1

# Meta Engine: A Unified Semantic Query Engine on Heterogeneous LLM-Based Query Systems

Abstract.
With the increasingly use of multi-modal data, semantic query has become more and more demanded in data management systems, which is an important way to access and analyze multi-modal data. As unstructured data, most information of multi-modal data (text, image, video, etc) hides in the semantics, which cannot be accessed by the traditional database queries like SQL. Comparing to traditional queries, semantic queries involve natural language to process the multi-modal data based on their semantics. Given the power of Large Language Model (LLM) in understanding semantics and processing natural language, in recent years several LLM-based semantic query systems have been proposed, to support semantic querying over unstructured data. However, this rapid growth has produced a fragmented ecosystem. Applications face significant integration challenges due to (1) disparate APIs of different semantic query systems and (2) a fundamental trade-off between specialization and generality. Many semantic query systems are highly specialized (e.g., DocETL for long documents, StructGPT for table QA), offering state-of-the-art performance within a single modality but struggling with multi-modal data. Conversely, some ”all-in-one” systems (e.g., LOTUS) handle multiple modalities but often exhibit suboptimal performance compared to their specialized counterparts in specific modalities. In short, the existing systems still have significant limitations when applied to complex multi-modal tasks.
This paper introduces Meta Engine, a novel ”query system on query systems”, designed to resolve those aforementioned challenges. Meta Engine is a unified semantic query engine that integrates heterogeneous, specialized LLM-based query systems. It combines the specialities of different systems to handle complex scenarios, and provides a unified natural language interface. Its architecture comprises five key components: (1) a Natural Language (NL) Query Parser to accept and break down complex NL queries into simpler sub-queries, (2) an Operator Generator that generates candidate operators for each sub-query, (3) a Query Router to dispatch sub-queries to the most appropriate specialized systems, (4) a set of Adapters to abstract and unify the heterogeneous APIs of underlying systems, and (5) a Result Aggregator to merge, deduplicate, and synthesize a coherent final answer. This design enables Meta Engine to combine the distinct strengths of different semantic query systems, achieving high-performance multi-modal semantic processing from a unified entry point. In the evaluation, Meta Engine consistently outperforms all baselines, yielding 3–6x higher F1 in most cases and up to  24x on specific datasets.
PVLDB Reference Format:
PVLDB, 14(1): XXX-XXX, 2020.
doi:$X$.XX/XXX.XX This work is licensed under the Creative Commons BY-NC-ND 4.0 International License. Visit https://creativecommons.org/licenses/by-nc-nd/4.0/ to view a copy of this license. For any use beyond those covered by this license, obtain permission by emailing info@vldb.org. Copyright is held by the owner/author(s). Publication rights licensed to the VLDB Endowment.
Proceedings of the VLDB Endowment, Vol. 14, No. 1 ISSN 2150-8097.
doi:$X$.XX/XXX.XX
**footnotetext: These authors contributed equally to this work.
PVLDB Artifact Availability:
The source code, data, and/or other artifacts have been made available at URL_TO_YOUR_ARTIFACTS.
1. Introduction
The advance of Large Language Models (LLMs) leads to a paradigm shift in data analytics, resulting in the development of advanced semantic query systems. (Cappuzzo et al., 2020; Toenshoff et al., 2023; Bandyopadhyay et al., 2020; Patel et al., 2024; Shankar et al., 2024; Jiang et al., 2023; Khattab et al., 2023; Zhang et al., 2025). They enable applications to semantically reason and analyze over complex (like multi-modal) datasets using natural language.
Despite this rapid progress, the current ecosystem of LLM-based query systems is highly fragmented, presenting critical challenges that hinder the comprehensive use of them on complex tasks. First, the different tools result in disparate and incompatible APIs. Usually each specific system requires a unique design due to their specific scenarios. Consequently, developers seeking to build applications that leverage the strengths of multiple systems are forced to write complex integration logic, increasing both development and maintenance overhead. Second challenge is a crucial trade-off between specialization and generality. Many state-of-the-art semantic query systems are highly specialized, achieving exceptional performance by optimizing for a single modality. For instance, DocETL (Shankar et al., 2024) excels at processing large-scale long text documents, while StructGPT (Jiang et al., 2023) is designed specifically for semantic querying of tabular data. Their high specialization in specific domain/modality prevents them from integrating and processing information across different modalities. Conversely, ”all-in-one” systems like LOTUS (Patel et al., 2024) attempt to provide a unified solution for multi-modal data (e.g., table, text, and image), but the cost of such a generalization is often suboptimal performance on specific single modality, like the image QA quality of LOTUS is usually lower than a pure VLM.
These challenges raise a critical need for a unified framework that can (1) comprehensively leverage the strengths of different specialized systems to achieve high-performance and truly multi-modal semantic query processing, and (2) provide a unified and easy-to-use interface for user interaction.
In this paper, we propose Meta Engine, a novel ”query system over query systems”. Meta Engine functions as an comprehensive framework, integrating heterogeneous LLM-based semantic query systems into a cohesive pipeline. By adaptively scheduling and calling the different systems, it can intelligently combine their specialities to achieve a significant performance gain. Additionally, Meta Engine provides a unified natural language (NL) querying interface, keeping the disparate APIs of underlying systems transparent to users. Note that although with an NL interface that allows users to query with NL questions, Meta Engine is not a question-answering (QA) system, but a general semantic query system. Meta Engine supports not only typical question answering tasks, but also broader data analytic tasks. In summary, the task of Meta Engine is modeled as such: given a supporting document repository with documents in different modalities (text, image and table), Meta Engine will process users’ queries and return the results. For example, in one of our evaluation dataset  (Rahman et al., 2025), each natural language query specifies an analytical task over structured tables, text and images. An example query and supporting documents are shown in Figure 1.
Figure 1. An End-to-End Multimodal Analytical Query Example in Meta Engine
Meta Engine includes several major components: (1) Natural Language Query Parser: This component accepts the user queries in natural language and intelligently preprocesses them. This mainly includes two modules, query complexity checker and query decomposer. The complexity checker determines whether the query is complex enough, and for a complex query (e.g., multi-hop queries), the decomposer will split it into a set of simpler sub-queries, each corresponding to a specific task and data modality, like the example shown in Figure 5.
Furthermore, we propose a novel hybrid sub-query generation method: initial decomposition + iterative sub-query refinement. Specifically, the query is first decomposed into sub-queries in one step, then the sub-queries are executed one by one while each sub-query is refined with the execution results from the proceeding sub-query if they have dependency. By such a dynamic refinement, the sub-queries are stated more clearly and the system avoids wasting time in duplicate computation across sub-queries.
(2) Operator Generator:
After decomposition, Meta Engine analyzes each sub-query’s intent and required evidence type, and maps it to a set of candidate operators using a confidence-based ranking mechanism rather than selecting a single operator. Specifically, the generator predicts a ranked list of operators, sorted by confidence scores, where higher-ranked operators are considered more likely to successfully answer the sub-query. To make routing efficient, this step relies only on lightweight metadata rather than full document contents.
(3) Query Router: The operators are then passed to the query router. The query router intelligently routes the opeartor to the most appropriate underlying specialist system (e.g., routing the text processing operator to DSPy and the image information extraction operator to VLM). Note that each operator could have multiple choices, like LOTUS, DSPy and VLM can all execute image operator. We design two types of routers: A learned router which is a query-dependent classifier on top of a pre-trained LLM, and a statistic router that is query-agnostic, based on history statistics. Our evaluation shows the learned router is overall better, but statistic router outperforms it in some cases.
(4) Adapters: Adapters serve as wrappers for the underlying LLM query systems. Each adapter conforms to a standardized internal interface while managing the unique API, request format, and authentication specifics of the system it wraps. This abstracts away the complexity and heterogeneity of the individual systems.
(5) Result Aggregator: As the adapters return results for their respective sub-queries, this final component aggregates the individual pieces of information and generates the final results. It refines the outputs (e.g., making long answers concise, extracting key answer keywords) and then merges and deduplicates (if necessary) the results from the different systems to synthesize a single, comprehensive, and accurate answer for the user.
By orchestrating the execution across specialized systems, Meta Engine resolves the specialist-generalist trade-off, enabling applications to benefit from high-performance, multi-modal processing through a single, unified natural language API.
The main contributions of this work are as follows:
  1. (1)
    We propose the Meta Engine architecture, a ”query system on query systems” that integrates diverse, specialized LLM query systems.
  2. (2)
    We introduce several components (Parser, Operator Generator, Router, Adapters, Aggregator) and novel optimization techniques to manage query decomposition, intelligent routing, execution and result synthesis, effectively combining the specialities of heterogeneous LLM query systems.
  3. (3)
    We integrate a unified natural language interface that abstracts the complexity of underlying APIs, simplifying the development of high-performance, multi-modal semantic applications on top of our system.
  4. (4)
    We conduct extensive evaluation that shows the effectiveness of our proposed system. We also perform ablation study to show the impacts of each major components in the system.
The remainder of this paper is structured as follows: Section 2 introduces prior works related to Meta Engine. Section 3 and 4 introduce the architecture, workflow and technical details of our method. Section 5 is the experiment section, analyzing the end-to-end results from different methods and the ablation studies for our method.
2. Related work
The study of semantic querying has historically aimed to bridge the gap between rigid data schemas and human-like understanding. In the pre-LLM era, visions like the Semantic Web (Lu et al., 2002; Spanos et al., 2012) embeds data semantics directly into the web via ontologies and knowledge graphs, but such approaches struggle with the high overhead of maintaining global ontologies. Subsequent approaches integrated advanced deep learning methods with databases. For example, (Cappuzzo et al., 2020; Toenshoff et al., 2023; Bandyopadhyay et al., 2020) utilize deep-learning-generated embeddings (e.g., word2vec (Mikolov et al., 2013)) and User-defined function (UDF) to compute semantic similarities between values or columns. However, these systems often lack a more comprehensive view of inter-column relationships and required users to manage complex UDFs. Recent work leverages large language models to perform high-level planning and optimization for semantic queries, automatically reasoning over query structure and execution strategies across complex data workflows (Wang et al., 2025). Furthermore, embedding is good at measuring the semantic similarity between objects, but cannot capture more complex semantics. For example, based on the embedding distance, users can determine how similar two values/objects are, but hard to know the relationship between them.
The advent of Large Language Models (LLMs) marked a paradigm shift, enabling systems to interpret flexible natural language predicates and execute complex reasoning without model modifications or extensive tuning. Recent semantic query systems focus on deeply integrating LLMs into data processing pipelines to enhance accuracy, usability and flexibility, benefiting from the strong semantic representation capabilities of large-scale language models across downstream  (Li et al., 2025). They usually define declarative programming interfaces (semantic operators) that extends the relational model with composable AI-based operators, powered by LLM. The operators will then be composed into complex semantic queries. For example, StructGPT (Jiang et al., 2023) represents a ”reading-then-reasoning” framework with operators like Extract_SubTable and Extract_Neighbor_Relations to process structured data using LLM. LOTUS (Patel et al., 2024) integrates semantic operators like semantic join and filter into Python Pandas Dataframe API, with performance optimization for each operator. UQE (Dai et al., 2024) integrates semantic operators into SQL in relational database environment, treating unstructured text as ”virtual columns” to allow seamless semantic queries in SQL. DocETL (Shankar et al., 2024) extends these capabilities to agentic document processing, designing LLM-powered operators to construct complex transformation pipelines for long-form documents. DSPy (Khattab et al., 2023) defines its declarative programming paradigm with a focus on prompt engineering, allowing developers to define high-level modules that the framework automatically optimizes the task pipeline for better prompts and weights. But those semantic query systems are either highly specialized (like DocETL focuses on long document processing) or general-purposed with average performance on most tasks.
3. Architecture
Figure 2. Architecture and End-to-End Execution Pipeline of Meta Engine for Multimodal Queries
Figure 2 presents the architecture and end-to-end workflow of Meta Engine to process a query.
3.1. Architecture
Meta Engine is designed as a modular system that coordinates the processing of complex, multimodal queries through a series of specialized components. The process initiates at the Query Parser (shown in Figure 2 Stage 1), which acts as the initial interface for user queries. This component functions as a complexity checker and query decomposer, inputting a query and outputting one or more sub-queries.
The Operator Generator (shown in Figure 2 Stage 2) is responsible for identifying and ranking the most appropriate analytical operators for each sub-query. It determines the optimal operator for each sub-query with the help of Query Router and Adapters.
The execution layer is comprised of the Query Router and various Adapters, such as those for LOTUS, DocETL, and DSPy which wrap existing semantic query systems. These adapters interface directly with the multi-modal data sources (tables, text, and images) to perform the required reasoning. Router takes charge of sending each sub-query to the proper adapter to execute.
Finally, the partial results from these isolated sub-query executions are synthesized by the Final Aggregator (shown in Figure 2 Stage 3). This component aggregates the answers from all sub-queries, filtering out noise and refining the output to produce a concise, accurate final answer for the user.
Meta Engine demonstrates a significant improvement in result quality compared to the state-of-the-art semantic query systems. By decomposing complex queries and routing them to specialized underlying systems, Meta Engine consistently achieves higher accuracy in our evaluation.
A core advantage of the architecture is its modular design, specifically the use of Adapters to wrap underlying LLM systems. This design decouples the reasoning logic from the backend execution, allowing for the seamless addition or removal of underlying systems (e.g., LOTUS, DSPy, DocETL) without overhauling the core engine. With new powerful systems emerge, they can be easily integrated as new adapters.
3.2. Workflow
The system will first check and potentially decompose the original user query, as shown in Figure 2 stage 1:
  1. (1)
    Query complexity checker determines whether the query is complex. If yes, go to step 2, otherwise, set the input query as a single sub-query and go to stage 2 directly.
  2. (2)
    Query decomposer generates initial sub-queries from the input query.
After we have all the sub-queries from stage 1, the system will generate and execute the operators from these sub-queries, as shown in Figure 2 stage 2:
  1. (1)
    For the current sub-query, Meta Engine generates a ranked list of possible operators to be processed with the data of related modality. The details are in Section 4.2.
  2. (2)
    The operators in the list will be routed to proper adapters in the order of rankings and executed one by one, until the first valid and meaningful result $R$ is returned from the execution. Then $R$ is the result for the current sub-query. Additionally, $R$ will be used to refine next sub-query, as introduced in Section 4.2.
  3. (3)
    Repeat Step 1 and 2 until all sub-queries are processed and their results $R_{1},R_{2},R_{3},$... are collected.
After we collected all sub-queries’ answers, the system will put everything into a final answer aggregator, as shown in Figure 2 stage 3, which will comprehend the results into the final result.
Such a workflow ensures (1) the complex query can be analyzed more accurately as single-hop sub-queries are more tractable by the LLM-based semantic systems; (2) the sub-queries are progressively refined with previous sub-query results, reducing sub-query ambiguity; (3) the final result is concise enough after the final aggregation. In summary, it effectively guarantees the result quality.
3.3. Operators
3.3.1. TextAnalytics
operator processes sub-queries whose supporting evidence resides in textual documents. Given a decomposed question and its associated passages, it extract relevant facts, entities, or relations and produces a modality-specific answer conditioned solely on the provided text.
3.3.2. TableAnalytics
operator is designed for processing sub-queries grounded in structured tabular data. It executes operations such as filtering, row/column selection, aggregation, comparison, and numerical computation according to the query intent through assigned adapters.
3.3.3. ImageAnalytics
operator answers sub-queries that require visual evidence from images, charts, or diagrams. It understands semantic information, including objects, labels, tr ends, and numeric values, and aligns these signals with the sub-query through multimodal reasoning.
4. Querying Stages and Optimization
In this section we introduce each components and the optimization on them.
Figure 3. Confidence-based Operator Ranking, Routing, and Execution in Meta Engine
4.1. Stage 1: Query Decomposer
You are a database query decomposer.
Your job is to decompose a complex question into
multiple simpler subqueries
that can be answered using the provided multimodal
data sources (table, image, text).
You must follow these formatting rules strictly:
1. The list must contain **at least one** and **at
   most three** subqueries.
2. Each subquery must be a short, clear, and self-
   contained QUESTION.
3. Maintain a logical basic-to-advance reasoning
   order (from basic retrieval to final inference).
We provide 3 types data about several entities
involved in the user query:
1. Table, that contains headers and several rows
2. Image, that either involves the entity or related
   to it.
3. Text, which include detailed description or
   relevent information about the entity.’
Few-Shots Examples
Figure 4. Prompt for Multimodal Query Decomposition
Figure 5. Example of Query Decomposition and operator generation in Meta Engine
4.1.1. Checker and Decomposer
Given a input query, Meta Engine first determines whether it should be answered directly or decomposed into several simpler sub-queries, for which a query complexity checker is invoked. The complexity checker determines whether the question can be answered in a single step or requires multi-step reasoning, with the help of LLM. Specifically, based on an LLM, the checker inspects whether the input query falls in any of the ”Complex Categories” below. If yes, the query is determined to be multi-hop.
  1. (1)
    Multi-part question: multiple tasks, multiple queries, explicit ”and/or/for each”.
  2. (2)
    Multi-step reasoning: needs combining at least 2 distinct facts or sub-results.
  3. (3)
    Aggregation or grouped analysis (sum/avg/count/trend/per-year/per-category).
  4. (4)
    Complex filtering (nontrivial constraints, conditions, temporal scopes).
  5. (5)
    Multi-entity comparison or ranking (compare X vs Y vs Z; find best among many).
  6. (6)
    Pipeline-style tasks (first do X, then do Y using X).
The checker also check whether the query is single-hop by looking for any ”Simple” signs:
  1. (1)
    One clear intent.
  2. (2)
    One entity or one fact.
  3. (3)
    One operator can answer it directly.
  4. (4)
    No decomposition improves clarity.
Particularly, if it is uncertain to the checker, the query is considered as multi-hop by default.
If the checker determines the query is single-hop, Meta Engine will bypass the decomposer and treats the original question as a single sub-query. When the query is classified as complex, the query decomposer module is triggered, with the prompt shown in Figure 4. The decomposer reformulates the original question into a set of simpler, self-contained sub-queries that are all single-hop. These sub-queries follow a basic-to-advanced reasoning order, where earlier sub-queries focus on retrieval of basic/grounding facts, and later sub-queries are built upon the results of preceding ones to support higher-level inference. As shown in Figure 5, the original query is decomposed into three sub-queries, where the first retrieves a basic fact about the flag while the last sub-query does advanced, higher-level analytics. Usually each sub-query can be answered within only one modality, e.g., based on one kind of data among text, images or tables.
4.1.2. Decomposer Optimization
To improve the overall system efficiency, we limit the number of sub-queries decomposed from each query to be at most three. This is because allowing unlimited decomposition may creates a significant number of sub-queries in some datasets. Those sub-queries are often unnecessarily detailed and there are often duplicates or large overlaps between them, which increases latency but does not actually produce better answers. Particularly, many of these sub-queries repeat similar reasoning or retrieve information that is not essential to answering the original query. In addition, long chains of sub-queries make the system more fragile. If an earlier sub-query produces an incorrect or incomplete result, that error will be passed to the later sub-queries and accumulated over time. This leads to slower end-to-end query processing and less accurate final result. By limiting the number of sub-queries, we encourage the system to focus on the most important analysis, reducing unnecessary computation and improving result stability for different datasets.
4.2. Stage 2: Sub-query Routing and Execution
type_of_operators:
imageAnalytics, tableAnalytics, textAnalytics.
You can output multiple possible operators for the
given question.
Each operator must include its type, question, data
sources, and confidence score (between 0 and 1).
Example outputs(multiple operators):
RANKING RUBRIC:
1. Direct Mention or Implication of Modality
- If the question explicitly mentions an image ...
  imageAnalytics highest.
- If the question explicitly mentions a table ...
  tableAnalytics highest.
- If the question asks for textual ... textAnalytics
  highest.
2. Nature of Information Requested
- Quantitative comparison ... tableAnalytics
  priority.
- Historical events ... textAnalytics priority.
- Appearance, colors ... imageAnalytics priority.
3. Ambiguous or Multi-modal Questions
- If multiple operator types could answer, ...
- If ambiguous, ...
4. Confidence Score Ranges (Recommended)
- Highest-ranked operator: 0.70–0.95
- Middle operator: 0.50–0.75
- Lowest operator: 0.30–0.60
The final answer MUST be strictly valid JSON and
follow this schema:
Figure 6. Prompt for Operator Generation and Confidence-Based Ranking
After sub-queries are produced, the next steps are operator generation and routing, which determines how each sub-query will be executed. Figure 5 shows an example on how a user query been decomposed and then further converted to operators. The workflow of operator generation and routing is shown in Figure 3. For each sub-query, the system performs the following steps:
  1. (1)
    Candidate operator generation: Given a sub-query, the operator generator proposes multiple candidate operators, each corresponding to a different modality (TextAnalytics, ImageAnalytics, or TableAnalytics). Instead of using full document contents, the generator relies only on document titles and short captions (e.g., one-sentence summaries for images), to keep routing efficient and low-cost. These candidates are generated by analyzing the intent of the sub-query together with these titles and captions as indicators of the available data sources. Each candidate operator is assigned a confidence score by LLM that estimates how suitable the modality is for answering the sub-query. The prompt is shown in Figure 6.
  2. (2)
    Confidence-based ranking and execution: The candidate operators are sorted in descending order of confidence. The system first executes the highest-ranked operator using its associated data sources. If the execution produces a valid answer, the process terminates and the result is returned. Otherwise, the system proceeds to the next operator in the ranked list. In some cases the top-ranked operator may not be able to properly process the sub-query, so such a progressively attempting is necessary.
  3. (3)
    Iterative fallback and refinement: If no valid answer is obtained by the higher-ranked operator, the system repeats the execution process with lower-ranked operators until an answer is found or all candidates have been exhausted. For sub-queries after the first one, the system additionally refines them before repeating the above steps. Specifically, next sub-query is refined by incorporating execution results from previous sub-queries, such as filling the resolved entity names or intermediate values from the results into next sub-query to reduce ambiguity.
Table 1. Statistics of the datasets
Dataset Queries Text Docs Tables Images Scenario
MultiModalQA 23,817 218,285 10,042 57,058 Multi-hop Q&A
ManyModalQA 1,440 1,440 1,440 1,440 Single-hop Q&A
Text2Vis 1,788 1,788 1,788 1,788 Analytical
FinMMR 1,291 1,291 0 2,269 Analytical
M2QA 1,000 1,884 4,686 37,972 Multi-hop Q&A
4.3. Stage 3: Aggregator and Final Result
You are an final answer aggregator. Read subqueries
and their tool outputs, then produce the best
possible answer to the original question using ONLY
the provided evidence. Do NOT invent facts.
Goal: Produce an EXTREMELY SHORT final answer based
on all provided Q&As.
Instructions:
1) Analyze all result fields.
2) When multiple answers are present:
    - If one answers the question and the other does
      not...
    - If both answer but differ...
    - If a later sub-query produces an answer that
      logically conflicts with an earlier one...
    - If one answer is descriptive and another...
3) If conflicts remain unresolved, prefer the answer
   that is more specific, complete, and contextually
   appropriate.
STRICT brevity rules for "final_answer":
- HARD CAP: <= 10 words total.
- If yes/no is possible: answer “Yes.” or “No.”
- Always give the most concise factual answer.
Figure 7. Prompt for Final Answer Aggregation with Conflict Resolution
After all sub-queries have been executed, the system put everything into a final answer aggregation module to produce a concise response to the original query from the sub-queries’ results. The inputs to this stage consist of a chain of sub-query–answer pairs, where each sub-query corresponds to an executed operator and its output. The aggregator analyzes all available sub-query results and selects those that directly supports the original query while filtering out irrelevant or non-informative outputs. The prompt is shown in Figure 7. When multiple candidate answers are present, the aggregator resolves conflicts between them by checking logical consistency across the reasoning chain and preferring answers that are coherent with previously established intermediate facts. The final output is aggregated under strict length constraints to encourage direct and unambiguous responses. By explicitly grounding the final answer in intermediate evidence and enforcing consistency across the reasoning chain, this aggregator completes the end-to-end pipeline with a stable and interpretable result.
4.4. Query Router
In this and the next section we introduce two components, the router and adapters, which are important but not discussed in previous sections in details.
For each operator, we utilized a router to select the most appropriate adapter to execute it. We construct two different types of routers, statistic and learned routers.
Statistic router is based on the statistics of best adapter for each operator over an annotated training set. Specifically, for the operator used in a training sub-query, we try all capable adapters (one by one) to execute the operator, and collect the performance metrics (e.g., Semantic Hit, introduced in Section 5.1) of each adapter’s execution. Then for real (testing) queries, each operator will be routed to the adapter with the best overall performance for it across all training sub-queries. In short, the statistic router will always route an operator to the best-performing adapter for it based on history results. We denote this statistic, static, sub-query agnostic router as StatRouter.
Learned router (denoted as MLRouter) is a learning-based, dynamic, sub-query dependent router. It accepts the sub-query, operator, and modality indicators as input, and learns to perform semantic-aware routing based on the sub-query and operator. We implement the learned router using Qwen-0.6B(Team, 2025) as a frozen encoder, followed by a lightweight multilayer perceptron (MLP) classifier. Only the parameters of the MLP are trained, while the backbone language model (Qwen) remains fully frozen throughout training and inference. Router training is conducted offline using historical execution traces collected from all the adapters over the training sub-queries. Specifically, for each sub-query and its operator, all capable adapters are executed individually, and their outputs are evaluated using multiple metrics, including F1, Hit, Coverage, and Semantic Hit (all introduced in Section 5.1). These metric scores are combined into a single quality score by weighted summation, where Semantic Hit and F1 scores are emphasized, while Hit and Coverage serve as secondary criteria. For each training sub-query $q$, the training label is the best adapter $k^{*}_{q}$ for executing this sub-query.
With such training data, the router can learn to choose the proper adapter dynamically based on the sub-query.
The training is performed using a weighted cross-entropy objective, where each training sub-query contributes proportionally to its sample weight. The sample weight for a training sub-query is computed as a function of (1) the absolute quality score of the best-performing adapter and (2) the margin between the best and sub-optimal adapters, as shown in Equation 1 and Equation 2 .
Formally, let $S_{q}^{k}\in[0,1]$ denote the aggregated quality score of adapter $k$ when executing sub-query $q$. Let $k^{*}_{q}=\arg\max_{k}S_{q}^{k}$ be the best-performing adapter, and define $S_{q}^{(1)}=\max_{k}S_{q}^{k}$ and $S_{q}^{(2)}$ as the second-best score among all adapters. Each training sub-query $q$ is associated with a confidence-aware sample weight $w_{q}$, defined as
(1) $w_{q}=\min\!\Big(w_{\max},\max\!\big(w_{\min},\alpha+\beta S_{q}^{(1)}+\gamma(S_{q}^{(1)}-S_{q}^{(2)})\big)\Big).$
The weight is bounded within $w_{\min},w_{\max}]$ to prevent overly dominant or vanishing contributions from individual training records. In our experiments, we set $\alpha=0.1$, $\beta=0.9$, $\gamma=0.2$, $w_{\min}=0.1$, and $w_{\max}=1.5$.
Given the router prediction $p_{\theta}(k\mid q)$ over adapters, training is performed using a weighted cross-entropy objective, to maximizes $p_{\theta}(k^{*}_{q}\mid q)$, i.e., the predicted probability of the groundtruth best adapter $k^{*}_{q}$ given sub-query $q$:
(2) $\mathcal{L}=\frac{1}{\sum_{q}w_{q}}\sum_{q}w_{q}\cdot\left(-\log p_{\theta}(k^{*}_{q}\mid q)\right).$
Such an objective considers not only the best adapter but also the sub-optimal adapters whose performance is close to the best one, which makes the router more robust, as the best adapter in training may not be still the best in testing, but it is likely still among the top adapters. The weighting mechanism allows the router to emphasize reliable routing signals without requiring hard filtering or manual thresholding of training data(Shazeer et al., 2017).
During inference, the router performs a single forward pass to produce a categorical adapter decision for each sub-query. As a result, both the statistic and learned routers raise negligible overhead of latency comparing to the downstream sub-query execution.
The routing module is fully decoupled from adapter implementations and downstream reasoning logic. New adapters can be added or removed by extending/shrinking the classification label space and retraining only the lightweight classifier, without modifying the backbone model or the rest of the system, which makes Meta Engine practical and adaptive to dynamic user demands.
4.5. Adapters
The adapters wrap the underlying semantic query systems (e.g., LOTUS, DSPy, etc) and provide unified APIs for executing the operators of Meta Engine. The unified interfaces allows transparently calling the diverse underlying systems.
Most of the adapters support all the three analytical operators in Meta Engine, i.e., TextAnalytics, TableAnalytics and ImageAnalytics, except StructGPT adapter which only supports TableAnalytics. Specifically, each operator has one interface (entry function) in each adapter which supports it. The interface accepts a sub-query and returns the execution result of the current operator over the sub-query.
Currently Meta Engine integrates four adapters, warping LOTUS, DSPy, Single-Model and StructGPT respectively. In this paper, we enforce deterministic decoding (temperature =0=0, fixed seed) to improve reproducibility and system stability.
LOTUS adapter supports all three modality operators. We implement each interface mainly using the semantic aggregation (sem_agg) in LOTUS. The supporting documents are stored in Pandas Dataframe by LOTUS, and sem_agg is called over them to comprehend the evidence of corresponding modality and analyze the sub-queries.
DSPy adapter implements each modality operator with DSPy, a programmatic prompting framework that treats each operator call as an explicit Language Model (LM) Program. DSPy factors the operator into reusable components (e.g., instruction template and constrained output format) and can compile these components into a stable program for repeated invocation across queries.
Single-model adapter implements the three modality operators via direct end-to-end prompting of a single vision-language model (VLM), serving as a strong baseline. For each operator, the adapter packages the supporting evidence in the corresponding modality and queries the VLM in one pass. In contrast to the other backend systems, the single-model adapter relies purely on the VLM’s ability to aggregate evidence and generate the final answer.
StructGPT adapter implements TableAnalytics operator with a structure-aware, multi-stage inference procedure that explicitly exploits the schema and row organization of the tables. For each sub-query, the adapter first performs schema grounding to select relevant columns for the sub-query, then it applies row filtering over the reduced schema, identifying candidate records that are likely to contain the answer. Finally, the adapter produces a response based on the candidate records.
The adapter framework is extensible: new backend systems can be easily integrated by wrapping into new adapters, providing the required interfaces and following the standard output format, without modifying the core execution pipeline. The router can also adapt to the new adapters with minimum cost, as discussed in Section 4.4.
Table 2. Main experiments results across 5 datasets. Metrics are reported as percentages (%).
Dataset Metric Lotus Single-Model LlamaIndex DSPy Meta-Engine (MLRouter) Meta-Engine (StatRouter)
Text2Vis F1-score 9.71 32.18 9.62 23.45 36.76 36.14
Hit 58.11 64.32 69.74 63.03 66.28 65.83
Semantic Hit 63.48 75.11 73.66 68.79 74.78 75.62
ManyModalQA F1-score 9.48 33.90 12.89 33.38 51.06 54.61
Hit 68.40 69.93 69.24 68.13 67.22 74.38
Semantic Hit 68.82 77.78 74.72 72.78 72.22 81.60
MultiModalQA F1-score 9.47 34.88 10.03 24.26 54.71 53.13
Hit 70.10 70.95 65.50 64.15 75.15 74.90
Semantic Hit 53.90 65.55 57.60 58.90 72.90 72.65
M2QA F1-score 5.43 26.67 10.14 28.27 47.91 50.57
Hit 46.11 43.51 77.25 77.05 65.07 69.86
Semantic Hit 38.72 39.12 62.08 64.87 62.87 66.47
FinMMR F1-score 1.16 20.75 2.92 19.70 26.90 28.47
Hit 11.62 25.41 42.37 25.02 35.40 37.41
Semantic Hit 16.42 34.86 53.99 34.24 50.43 52.59
5. Experiments
5.1. Experiment settings
All experiments are evaluated on a machine with one NVIDIA B200 80GB Tensor Core GPU, 3 AMD Rome cores CPU and 20GB RAM.
5.1.1. Datasets
We evaluate Meta Engine on several publicly available multi-modal datasets: MultiModalQA (Talmor et al., 2021), ManyModalQA (Hannan et al., 2020), Text2Vis (Rahman et al., 2025), M2QA  (Abaskohi et al., 2025) and FinMMR (Tang et al., 2025). Table 1 summarizes their statistics. All datasets involve supporting data from at least two of the three modalities: text, tables, and images. Each query in those datasets is supposed to be answered based on the supporting textual, tabular, and/or visual information. And we unify the supporting documents from all datasets into a standardized multimodal document repository.
ManyModalQA is a question-answering (QA) dataset including questions, supporting data from the three modalities above, and ground-truth answers, where many questions are single-hop.
MultiModalQA is another QA dataset including questions, answers and multimodal supporting data, too. But its questions are usually multi-hop.
Text2Vis is a multi-modal analytics dataset, where each record includes a natural language query for a analytic task, as well as supporting documents: data tables, background introduction about the data, and charts based on the tables. Particularly, the charts are generated by ourselves. The original Text2Vis dataset includes Python code for visualizing the key calculation/statistics based on the table and query. We run the code to generate the charts as supporting images.
FinMMR is a multimodal benchmark focusing on financial numerical reasoning, where each query requires integrating financial knowledge with visual evidence such as tables, charts, and structured diagrams. All questions demand precise numerical results.
M2QA is a multimodal question answering dataset constructed from real-world Wikipedia articles, where each record is associated with a question and a source Wikipedia URL, where the question needs to be answered based on the Wiki page. To use this dataset, we retrieve and separate different modality data from the referenced Wiki pages. The extracted text, images, and tables are then normalized into our unified multimodal document repository. In the evaluation on this dataset, we skip the queries whose supporting images are not fully valid, caused by retrieval issues in preprocessing. Finally there are 501 queries used in the evaluation.
Note that in all experiments, due to the non-trivial latency of LLM-based systems, we use subsets of each dataset rather than their full collections. For MultiModalQA we randomly sample 2000 queries from the 23817 queries, while for ManyModalQA, Text2Vis, and FinMMR, we use the whole datasets. For M2QA, as mentioned above, we use 501 queries.
5.1.2. Methods
We evaluate our approach against several representative baselines: Single-Model, DSPy (Khattab et al., 2023), LOTUS (Patel et al., 2024), and LlamaIndex (Liu, 2022). All baselines follow a unified implementation pattern for multiple modalities. Specifically, since the baseline systems like LOTUS and DSPy usually do not well handle cross-modal interactions (like processing text and image together for one query), we implement the baseline methods by separating-then-aggregating: Each input query is processed within each single modality of the three, i.e., analyzing it with supporting text, images and tables separately. Then the results from the three modalities will be comprehended by an LLM into the final result. All methods, including Meta Engine, are using GPT-4.1-mini as the backbone LLM in this paper. Implementations for each baseline are shown below:
Single Model baseline is a single multi-modal LLM, i.e., GPT-4.1-mini in this paper.
LOTUS is a Pandas (pandas development team, 2020) DataFrame based semantic query system, with semantic operators powered by LLM. In this baseline, we mainly use LOTUS’s semantic aggregation operator (sem_aggsem\_agg) for processing each modality and final aggregation. This baseline will aggregate the supporting documents in each modality, get the results, and aggregate the results into the final result.
DSPY is a programmatic prompting framework that lets users define AI pipelines as modular programs (signatures + modules) and execute them with an underlying LLM/VLM backend. In this baseline,
we instantiate a ChainOfThought module to analyze the original question separately for each modality by conditioning only on the corresponding supporting evidence.
Then, we use a lightweight Predict module as an aggregator to generate the final response.
LlamaIndex is a programming framework for building LLM-based workflows over heterogeneous data sources. In this baseline, we use its indexing and querying abstractions to implement a retrieval-augmented generation (RAG) workflow for text and table modalities, and a direct vision-language querying workflow for images.
The final answer is generated by synthesizing the modality-level outputs with an LLM.
Meta-Engine, as we mentioned in the previous parts,
we propose two variants under the same framework: Meta-Engine (MLRouter) and Meta-Engine (StatRouter). Both use identical components and configurations for complexity checking, decomposition, execution and aggregation, differing only in the routing strategy. MLRouter is the learned router, while StatRouter is the statistic router.
5.1.3. Metrics
We evaluate end-to-end answer quality using both token-based metrics and semantic similarity metrics. For datasets that provide multiple acceptable ground-truth answers per question, we compute each metric between the prediction and every ground-truth answer, and report the maximum score to account for answer variability.
Let $\hat{y}$ denote the predicted answer and $y$ is a ground-truth answer. After normalization (lowercasing, punctuation removal, and whitespace trimming), we tokenize both answers into token multisets (i.e., bags of tokens preserving multiplicity). Let T(y^)T($\hat${y}) and $T(y)$ denote the resulting multisets, and let
(3) $m(\hat{y},y)=|T(\hat{y})\cap T(y)|$
be the multiset overlap count.
Token-level F1.
We compute precision and recall from token overlap:
(4) $\mathrm{Prec}(\hat{y},y)=\frac{m(\hat{y},y)}{|T(\hat{y})|},\qquad\mathrm{Rec}(\hat{y},y)=\frac{m(\hat{y},y)}{|T(y)|},$
and the token-level F1 score:
(5) $\mathrm{F1}(\hat{y},y)=\frac{2\cdot\mathrm{Prec}\cdot\mathrm{Rec}}{\mathrm{Prec}+\mathrm{Rec}}.$
Hit and Semantic Hit.
To complement strict lexical matching, we report two overlap-based metrics. Hit is a binary metric indicating whether $\hat{y}$ shares at least one token with $y$:
(6) $\mathrm{Hit}(\hat{y},y)=\mathbb{I}[m(\hat{y},y)>0].$
However, token-overlap metrics may penalize valid paraphrases or fail to capture subtle factual errors. To obtain a higher-level semantic assessment, we additionally employ an LLM as a semantic hit judge. Given the question, the prediction $\hat{y}$, and the set of ground-truth answers to the question, the judge determines whether $\hat{y}$ conveys the same factual content with at least one ground-truth answer, ignoring surface-form variations such as paraphrasing or reordering. If yes, $\hat{y}$ is a semantically correct answer, and its Semantic Hit is 1, as shown below:
(7) $\mathrm{Sem\_Hit}(\hat{y},y)=\mathbb{I}[\hat{y}\ is\ semantically\ equivalent\ to\ y].$
In the evaluation results, we report the average Hit and average Semantic Hit, i.e., the fraction of hit questions over all questions, and the fraction of semantically hit questions over all.
In addition, we also measure the end-to-end query processing time as efficiency metric.
5.2. End-to-End Evaluation
Across all datasets, Meta Engine achieves the strongest end-to-end result quality in most cases, showing improvements over all evaluated baselines across a wide range of scenarios, particularly on queries requiring multi-step reasoning and cross-modal evidence integration like MultiModalQA.
In terms of both F1-score and semantic correctness, Meta Engine exhibits clear and substantial advantages. These results indicated that advantages are not confined to a single dataset or modality, but generally across all datasets. In a few cases, LlamaIndex achieves higher Hit scores, but its Semantic Hit and F1-score are consistently lower than that of Meta Engine. This suggests that LlamaIndex tends to produce long, verbose answers which are incorrect semantically, making no sense, while Meta Engine delivers more semantically accurate results.
Specifically, according to Table 2, Meta Engine shows clear gains over the strongest competing baselines on most datasets. On Text2Vis, Meta Engine (MLRouter) achieves an F1-score of 36.76, exceeding the best baseline (Single-Model, 32.18) by 4.6%. On ManyModalQA, Meta Engine (StatRouter) reaches an F1-score of 54.61, outperforming the strongest baseline by 20.7%, while improving Semantic Hit from 77.78 to 81.60. On MultiModalQA, Meta Engine (MLRouter) attains an F1-score of 54.71, surpassing the best baseline by 19.8%, together with a Semantic Hit increase of over 7%. On M2QA, Meta Engine (StatRouter) achieves an F1-score of 50.57, improving upon the strongest baseline by 22.3%, and yields the highest Semantic Hit (66.47). On FinMMR, Meta Engine attains an F1-score of 28.47, which is comparable to LlamaIndex, while remaining competitive in terms of semantic correctness across all baselines.
Existing baselines typically attempt to answer complex queries in a single pass or through a fixed pipeline, forcing a single model or framework to handle heterogeneous reasoning requirements simultaneously. In contrast, Meta Engine decomposes complex queries into simpler, modality-aligned sub-queries, ensuring that each reasoning step is grounded in the most appropriate data modality and reducing cross-modal interference during reasoning. Also, Meta Engine overcomes this limitation by dynamically routing each sub-query to the most appropriate adapter, effectively combining the strengths of multiple specialized systems. This coordination mechanism enables Meta Engine to leverage modality-specific expertise during execution, leading to consistently stronger overall performance.
While some baselines achieve moderately higher hit rates through more token overlap with ground-truth answers, they often produce incomplete, verbose, or weakly grounded responses. In contrast, Meta Engine consistently attains higher semantic accuracy, indicating that its answers better capture the intended factual content of the queries. These improvements come from the effective Meta Engine architecture of query decomposition, operator generation, routing, and multi-adapter execution.
Furthermore, we apply a lightweight post-processing step for the baselines which tends to output verbose answers (such as LOTUS), to removing the leading explanatory text and extracts the major content from their answers. However, even after this step, the resulting baseline answers still remain long, reflected on the low F1 scores of LOTUS and LlamaIndex. This proves that simple addition to the existing semantic query systems cannot effectively improve them, which strengthens the necessity of the new architecture and workflow in Meta Engine.
Table 3. Ablation results for MultiModalQA. All metrics are reported as percentages (%).
MultiModalQA (len = 2000)
Method F1 (%) Hit (%) Semantic Hit (%)
MLRouter
Meta-Engine 54.71 75.15 72.90
Meta-Engine w/o Aggregator 52.75 74.15 70.60
Meta-Engine w/o Decomposer 54.66 72.70 71.40
Meta-Engine w/o Checker 52.60 76.10 71.95
StatRouter
Meta-Engine 53.13 74.90 72.65
Meta-Engine w/o Aggregator 40.79 75.15 71.60
Meta-Engine w/o Decomposer 53.01 73.05 71.00
Meta-Engine w/o Checker 53.18 76.80 73.85
Single-Adapter (LOTUS)
Meta-Engine 39.76 62.35 57.15
Table 4. Ablation results for ManyModalQA. All metrics are reported as percentages (%).
ManyModalQA (len = 1440)
Method F1 (%) Hit (%) Semantic Hit (%)
MLRouter
Meta-Engine 51.06 67.22 72.22
Meta-Engine w/o Aggregator 50.50 68.19 72.50
Meta-Engine w/o Decomposer 51.47 68.19 72.64
Meta-Engine w/o Checker 51.09 72.78 76.67
StatRouter
Meta-Engine 54.61 74.38 81.60
Meta-Engine w/o Aggregator 46.80 74.86 81.25
Meta-Engine w/o Decomposer 54.40 74.38 80.14
Meta-Engine w/o Checker 51.00 73.47 78.89
Single-Adapter (LOTUS)
Meta-Engine 34.83 53.54 57.85
Table 5. Ablation results for M2QA. All metrics are reported as percentages (%).
M2QA (len = 501)
Method F1 (%) Hit (%) Semantic Hit (%)
MLRouter
Meta-Engine 47.91 65.07 62.87
Meta-Engine w/o Aggregator 35.62 59.08 52.50
Meta-Engine w/o Decomposer 47.47 60.28 58.48
Meta-Engine w/o Checker 43.88 65.47 61.28
StatRouter
Meta-Engine 50.57 69.86 66.47
Meta-Engine w/o Aggregator 29.20 66.27 57.29
Meta-Engine w/o Decomposer 51.02 68.86 65.47
Meta-Engine w/o Checker 45.80 67.86 64.67
Single-Adapter (LOTUS)
Meta-Engine 46.58 69.06 65.27
5.3. Ablation Study
To understand the contribution of individual components in the Meta-Engine, we conduct a series of ablation experiments by selectively disabling specific modules. Each ablation isolates a design choice and reveals its functional role in the overall system. The results are shown in Table 3, 4 and 5.
5.3.1. Without Checker
In this setting, the complexity checker is disabled, and all queries are treated as complex by default. As a result, every input query is decomposed into sub-queries, regardless of whether decomposition is actually necessary.
Without checker, the result quality is usually lower in most cases. Although there are a few exceptions (Table 3), this does not affect the necessity of the checker, as checker improves the end-to-end efficiency significantly. As in Table 6, decomposing all queries without the checker could raise up to 70% more processing time. By preventing decomposing simple queries, a non-trivial portion of queries are processed in only one pass.
5.3.2. Without Decomposer
In the w/o Decomposer setting, the system never performs query decomposition. All original questions are considered non-complex and are passed directly into the pipeline as a single sub-query.
Removing decomposer makes the system overall perform worse, proving the effectiveness of query decomposition. There is only one exception in Table 4. This is because ManyModalQA has many single-hop queries, so decomposing everything may raises extra noise.
5.3.3. Without Aggregator
In this experiment, the final answer aggregator is disabled. While the system still decomposes the query and executes sub-queries iteratively, the final output is taken directly from the last sub-query’s answer, rather than being aggregating from all intermediate results.
In all cases the result quality drops significantly, especially F1-score score and Semantic Hit. This is because our aggregator comprehend the sub-queries results and ensure the final result is concise and more accurate. Hit score may be better without aggregator, but this is simply because the system tends to output long result, which has a larger chance to hit the ground-truth keywords but is actually not a correct answer.
5.3.4. Single Adapter (LOTUS)
We also evaluate the impact of the number of adapters. In the Single Adapter (LOTUS) configuration (the last row in the three tables), the system uses only the LOTUS adapter and disables the router as there is only one adapter to route. All other components from checker to aggregator are used as normal.
Comparing to the full versions of Meta Engine, single adapter system shows a lower performance, but it still outperforms the corresponding baseline method. Specifically, single LOTUS adapter Meta Engine achieves much higher F1-score (39.76 vs 9.47) and Semantic Hit (57.14 vs 53.90) on MultiModalQA, and all three metrics are significantly better than the LOTUS baseline on M2QA: F1-score 46.58 vs 5.43, Hit 69.06 vs 46.11, Semantic Hit 65.27 vs 38.72. These mean that (1) more adapters do increase the overall performance of Meta Engine, proving that Meta Engine effectively combines the specialities from different semantic systems; (2) The architecture and workflow design of Meta Engine effectively enhances the query processing quality of the underlying semantic systems.
Table 6. Average Query Runtime comparison (in seconds) including all ablations.
Method ManyModalQA MultiModalQA M2QA
Lotus 6.89 7.55 6.13
Single-Model 7.08 8.01 6.87
LlamaIndex 4.15 5.99 4.81
DSPy 7.37 9.03 11.63
Meta-Engine (MLRouter)
Full 8.15 11.91 19.89
w/o Aggregator 6.45 11.76 17.00
w/o Decomposer 5.76 7.33 8.30
w/o Checker 13.04 19.26 23.58
Meta-Engine (StatRouter)
Full 8.72 13.08 21.01
w/o Aggregator 7.53 12.26 20.99
w/o Decomposer 6.12 9.02 10.46
w/o Checker 14.53 18.74 25.81
5.4. Efficiency Analysis
We measure end-to-end efficiency on some datasets and report in Table 6. The result shows that Meta Engine achieves competitive runtime across most datasets. On ManyModalQA and MultiModalQA, most Meta Engine variants show a similar or faster running speed than baselines like DSPy and Single-Model, and remain close to lightweight baselines such as LlamaIndex. Meta Engine without checker, as we mentioned in Section 5.3, usually raises the latency, so it has the longest runtime. Although the full version of Meta Engine is slightly slower in some cases like the StatRouter version on MultiModalQA, the overhead remains moderate.
We will further improve the efficiency in our future work. Specifically, we observe that the latency is mainly raised in the operator ranking and generating stage, and we also observe this is depending on the dataset:
the operator generation attempts averagely 1.05 - 1.13 iterations per query on MultiModalQA, wihle it attempts 1.16 - 1.28 on ManyModalQA, and 1.52 - 1.81 on M2QA, which is positively correlated with the runtime on these datasets. This indicates the operator ranking is a key to guaranteeing the efficiency. So we will improve its accuracy to minimize the iterations in our future work. In addition, we plan to further improve efficiency through parallel sub-query processing: some sub-queries may have no dependency between each other, so they can be processed parallely instead of sequentially, which will effectively resolve the efficiency bottleneck. We will also apply more optimizations such as parallel execution of multiple queries and result caching for similar queries or sub-queries.
6. Conclusion
In this paper, we presented Meta Engine, a unified semantic query engine that coordinates multiple LLM-based query systems to support complex multimodal tasks. Instead of relying on a single monolithic model or a fixed execution pipeline, Meta Engine decomposes complex queries into modality-aligned sub-queries, routes each sub-query to the most suitable specialized system, and aggregates intermediate results to produce accurate and coherent final answers.
Extensive experimental results across diverse multimodal benchmarks demonstrate that Meta Engine achieves superior end-to-end answer quality in most settings, particularly on tasks requiring multi-step reasoning and multi-hop evidence integration.
Overall, this work presents a system that explicitly decomposes queries and coordinates multiple specialized systems. By enabling flexible integration and coordination of LLM-based specialized systems, Meta Engine provides a practical and extensible foundation for building high-performance multimodal semantic query applications.